Carbon dioxide produced by yeast in bread dough causes the dough to rise before baking during the baking process The carbon dioxide expands predict the final volume of 0.1 L carbon dioxide gas trapped in the bread dough as it’s heat it from 25°C to 98°C
1
Expert's answer
2020-11-20T10:05:29-0500

"\begin{aligned}
V_1 &= 0.1L\\
T_1 &= 25°C = (273 + 25)\ 298K\\
T_2 &= 98°C = (273 + 98)\ 371K\\
V_2 &=\ ?
\end{aligned}"


According to Charles's law,


"\dfrac{V_1}{T_1} = \dfrac{V_2}{T_2}"


"\dfrac{0.1}{298} = \dfrac{V_2}{395}"


"V_2 = \dfrac{395×0.1}{298}"


"V_2 = 0.133L"


"\therefore" The final volume of the carbon dioxide gas is 0.133L.